OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 3x2 − 3x − 60

OpenStudy (anonymous):

1. you need two factors, whose sum is equivalent to the middle term of the expression (-3x) and whose product is equal to the the product of first and last terms (3x^2 * -60) you follow upto there?

OpenStudy (anonymous):

it might be easier if you rewrite the problem -3x + 3*2 - 60 you want to pull out the common factor (3) 3(-x + 2 - 20) or -3(x -2 + 20) does this answer it or do i need to break it down more?

OpenStudy (anonymous):

a little more

OpenStudy (campbell_st):

take 3 out as a common factor \[3(x^2 - x - 20) \] next find the factors of 20 that add to -1.... the larger factor is negative... -20 x 1 -20 + 1 = -19 -10 x 2 -10 + 2 = -8 - 5 x 4 -5 + 4 = -1 the factors are -5 and 4 then its \[3(x - 5)(x + 4) = 3x^2 - 3x - 60\]

OpenStudy (anonymous):

2. those factors are 12x and -15x (note -15x+12x = -3x and -15x * 12x = -180x^2) so now substitute this two factors for the middle term and your expression now becomes: \[3x ^{2} +12x - 15x -60\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

divide the whole equation by three (3(-x + 2 - 20))/3 this will give you -x + 2/3 - 20/3 = -x -18/3 or -1(x + 18/3) ...it has been a while since i've done math but i didn't know if in your original equation if that was supposed to be an "x" or a times symbol

OpenStudy (anonymous):

@simplybeyoutiful \[3x(x+4)-15(x+4)\] =\[(3x-15)(x+4)\] that's how to completely factor